Use the following data to answer questions in this part:

Balance sheet data

Assets

20X7

20X6

Cash

$2,900

$1,000

Accounts receivable

2,500

2,000

Inventory

7,400

8,000

Property, plant, equipment

9,200

9,000

Accumulated depreciation

(2,900)

(2,500)

Total assets

$19,100

$17,500

Liabilities and Equity

Accounts payable

$4,700

$4,500

Interest payable

1,500

1,000

Dividends payable

1,000

2,500

Long-term debt

4,350

3,700

Bank note

1,000

800

Common stock

3,300

3,000

Retained earnings

3,250

2,000

Total liabilities and equity

$19,100

$17,500

Income statement for the year 20X7

Sales

$28,500

COGS

19,900

Depreciation

3,700

Interest expense

1,244

Gain on sale of old machine

1,150

Taxes

1,142

Net income

$3,664

Notes:

  • Dividends declared to shareholders were $850.
  • New common shares were sold at par for $1020.
  • Fixed assets were sold for $3500. Original cost of these assets was $2000, and $350 of accumulated depreciation has been charged to their original cost.
  • The firm borrowed $200 in a 10-year bank note – the proceeds of the loan were used to pay for new fixed assets.
  • Depreciation for the year was $3700 (accumulated depreciation up $2900 and depreciation on sold assets $800).
  • The company uses the LIFO inventory cost flow method. Had FIFO been used, inventories would have been $1,000 higher in 20X6 and $900 higher in 20X5.
  • The effective tax rate for 20X6 was 30%. For all other years, the effective tax rate was 20%.
  • At the beginning of 20X8, A firm issues a $10,000 bond with a 6% coupon rate, 4-year maturity, and annual interest payments when market interest rates are 7%.

    Use the following data to answer questions in this part:

    Balance sheet data

    Assets

    20X7

    20X6

    Cash

    $2,900

    $1,000

    Accounts receivable

    2,500

    2,000

    Inventory

    7,400

    8,000

    Property, plant, equipment

    9,200

    9,000

    Accumulated depreciation

    (2,900)

    (2,500)

    Total assets

    $19,100

    $17,500

    Liabilities and Equity

    Accounts payable

    $4,700

    $4,500

    Interest payable

    1,500

    1,000

    Dividends payable

    1,000

    2,500

    Long-term debt

    4,350

    3,700

    Bank note

    1,000

    800

    Common stock

    3,300

    3,000

    Retained earnings

    3,250

    2,000

    Total liabilities and equity

    $19,100

    $17,500

    Income statement for the year 20X7

    Sales

    $28,500

    COGS

    19,900

    Depreciation

    3,700

    Interest expense

    1,244

    Gain on sale of old machine

    1,150

    Taxes

    1,142

    Net income

    $3,664

    Notes:

  • Dividends declared to shareholders were $850.
  • New common shares were sold at par for $1020.
  • Fixed assets were sold for $3500. Original cost of these assets was $2000, and $350 of accumulated depreciation has been charged to their original cost.
  • The firm borrowed $200 in a 10-year bank note – the proceeds of the loan were used to pay for new fixed assets.
  • Depreciation for the year was $3700 (accumulated depreciation up $2900 and depreciation on sold assets $800).
  • The company uses the LIFO inventory cost flow method. Had FIFO been used, inventories would have been $1,000 higher in 20X6 and $900 higher in 20X5.
  • The effective tax rate for 20X6 was 30%. For all other years, the effective tax rate was 20%.
  • At the beginning of 20X8, A firm issues a $10,000 bond with a 6% coupon rate, 4-year maturity, and annual interest payments when market interest rates are 7%.
  • Calculate the current ratio for 20X6 for both LIFO and FIFO inventory cost flow methods. Consider all payables as current liabilities. Please explain each step of your calculation and interpret your results?
Answer #1

CALCULATION OF CURRENT RATIO CONSIDERING LIFO INVENTORY COST FLOW METHODS FOR 20X6

CURRENT RATIO = CURRENT ASSETS/CURRENT LIABILITIES

CURRENT ASSETS = CASH + ACCOUNTS RECEIVABLES + INVENTORY = 1000 + 2000 + 8000 = $11,000

CURRENT LIABILITIES = ACCOUNTS PAYABLE + INTEREST PAYABLE + DIVIDENDS PAYABLE = 4500 + 1000 + 2500 = $8,000

CURRENT RATIO = 11,000/8,000 = 1.375

CALCULATION OF CURRENT RATIO CONSIDERING FIFO INVENTORY COST FLOW METHODS FOR 20X6

CURRENT RATIO = CURRENT ASSETS/CURRENT LIABILITIES

INVENTORY AS PER FIFO METHOD => $1,000 HIGHER THAN INVENTORIES AS PER LIFO METHOD, THEREFORE INVENTORIES AS PER FIFO METHOD WOULD BE (INVENTORIES AS PER LIFO METHOD + $1,000) = 8000 + 1000 = $9,000

CURRENT ASSETS = CASH + ACCOUNTS RECEIVABLES + INVENTORY = 1000 + 2000 + 9000 = $12,000

CURRENT LIABILITIES = ACCOUNTS PAYABLE + INTEREST PAYABLE + DIVIDENDS PAYABLE = 4500 + 1000 + 2500 = $8,000

CURRENT RATIO = 12,000/8,000 = 1.5

ANALYSIS:

IN MOST INDUSTRIES, CURRENT RATIO OF ATLEAST 1.2 IS CONSIDERED APPROPRIATE. IN THE GIVEN CASE, WHEN CURRENT IS RATIO CALCULATED (USING EITHER CASES OF INVENTORY VALUATION) IN THE GIVEN QUESTION, IT TURNS OUT TO BE GREATER THAN 1.2. THEREFORE, IT IS TO BE SAID THAT COMPANY IS HAVING GOOD RATIO.